OpenStudy (anonymous):

I need help An Ice Skater starts with a velocity of 2.25 in 50.0 degree direction. After 8.33s, she is moving 4.65m/s in 120 degree direction. What are the X and Y components of her acceleration.

OpenStudy (anonymous):

 Solving the trig I think the answer is -0.32 m/s/s in X direction and + 0.64 m/s/s in Y direction.

OpenStudy (theeric):

Is this using\[\Large a_x=\frac{v_{f,x} - v_{i,x}}{t_{f,x} - t_{i,x}}\]?

OpenStudy (theeric):

And splitting up the components like

OpenStudy (anonymous):

I am assuming 50 and 120 angles w.r.t an arbitrary line which i am calling x-axis Split components like this Initial velocity: vix = 2.25 cos50 viy=2.25 sin50 Final velocity vfx= 4.65 cos120 vfy =4.65 sin120 Then find components of acceleration in both x and y directions ax = vfx-vix / t ay = vfy-viy / t t=8.33s=time taken for the change in the velocity.